# The Instability of Safety: How Random Seeds and Temperature Expose Inconsistent LLM Refusal Behavior

### v1

@@ Abstract
Current safety evaluations of large language models rely on single-shot testing, implicitly assuming that model responses are deterministic and representative of the model’s safety alignment. We challenge this assumption by investigating the stability of safety refusal decisions across random seeds and temperature settings. Testing four instruction-tuned models from three families (Llama 3.1 8B, Qwen 2.5 7B, Qwen 3 8B, Gemma 3 12B) on 876 harmful prompts across 20 different sampling configurations (4 temperatures $\times$ 5 random seeds), we find that 18–28% of prompts exhibit decision flips—the model refuses in some configurations but complies in others—depending on the model. Our Safety Stability Index (SSI) reveals that higher temperatures significantly reduce decision stability ($\chi^{2}=133.15$, $p<0.001$), with mean SSI dropping from 0.951 at temperature 0.0 to 0.896 at temperature 1.0. We validate our findings across all model families using Claude 3.5 Haiku as a unified external judge, achieving 91.7% inter-judge agreement with our primary Llama 70B judge (Cohen’s $\kappa=0.744$). The most stable model (Gemma 3 12B, SSI=0.965) is also the least restrictive (78.5% refusal rate), suggesting a potential stability-conservatism tradeoff. These findings demonstrate that single-shot safety evaluations are insufficient for reliable safety assessment and that evaluation protocols must account for stochastic variation in model behavior. We show that single-shot evaluation agrees with multi-sample ground truth only 92.4% of the time, and recommend using at least 3 samples per prompt for reliable safety assessment.
@@ 1 Introduction
  • Temperature significantly affects stability ($\chi^{2}=133.15$, $p<0.001$): Lower temperatures yield more stable decisions (mean SSI = 0.951 at temp 0.0) compared to higher temperatures (mean SSI = 0.896 at temp 1.0)
@@ 3.1 Experimental Design
Dataset. We use 876 harmful prompts from the BeaverTails dataset [6], covering categories including violence, hate speech, illegal activities, and other harmful content.
@@ 4.1 Overall Stability
The overall response distribution shows that the model predominantly refuses harmful requests (83.1% REFUSE, 10.9% PARTIAL, 5.9% COMPLY), suggesting strong safety alignment on average. However, the substantial proportion of unstable prompts indicates that aggregate statistics obscure significant per-prompt variability.
@@ 4.2 Distribution of Stability Scores
Figure 2 presents the distribution of SSI scores across all 876 prompts. The distribution is heavily right-skewed with a strong mode at perfect stability (SSI = 1.0): the median SSI is 1.0, indicating that more than half of prompts produce identical responses across all configurations. However, 14.3% of prompts fall below the 0.8 stability threshold, demonstrating that instability affects a meaningful fraction of safety evaluations.
@@ 4.3 Temperature Effects
Figure 3 illustrates the relationship between temperature and both compliance rate and mean SSI. We observe that stability decreases monotonically as temperature increases, with this effect being highly significant ($\chi^{2}=133.15$, $p<0.001$; Friedman $\chi^{2}=44.71$, $p<0.001$). The Friedman test is appropriate here because the same prompts are tested at each temperature (repeated measures design). Mean SSI drops from 0.951 at temperature 0.0 to 0.896 at temperature 1.0. Post-hoc Wilcoxon signed-rank tests with Bonferroni correction confirm significant decreases in SSI between temperature 0.0 and 1.0 ($p<0.001$). The flip rate more than doubles from 9.5% at temperature 0.0 to 19.6% at temperature 1.0.
@@
Interestingly, the lowest compliance rate occurs at temperature 1.0 (3.0%). However, this likely reflects that high-temperature sampling produces less coherent outputs that fail to meaningfully engage with the harmful request, rather than representing more robust safety behavior.
@@ 4.5 What Predicts Prompt Instability?
To understand what makes certain prompts unstable, we computed linguistic features for each prompt and analyzed their correlation with SSI. We find that prompt characteristics meaningfully predict instability:
  • Sensitive terms increase stability: Prompts containing obvious harmful keywords (e.g., “hack,” “bomb,” “kill”) have higher SSI ($r=0.18$), suggesting the model confidently refuses clearly harmful requests.
  • Multi-sentence prompts are less stable: Prompts with multiple sentences show lower SSI (r=0.21r=-0.21), possibly because complex requests create ambiguity.
A logistic regression classifier predicting instability from linguistic features achieves AUC = 0.70 in 5-fold cross-validation, indicating that surface-level features partially explain which prompts will be unstable. The most predictive features are average word length, sensitive term count, and word count.
Qualitatively, the most unstable prompts (SSI << 0.5) are predominantly copyright-related requests (e.g., requests for song lyrics or book passages) where the model oscillates between refusing, providing partial content with disclaimers, and complying. This suggests copyright represents a particularly ambiguous category in the model’s safety training.
@@ 4.6 Model Comparison
Figure 5 compares stability across the four evaluated models. We observe consistent patterns of instability across all three model families, with instability rates (SSI << 0.8) ranging from 14% to 28%.
@@
Table 2: Stability comparison across model families. All models evaluated using Claude 3.5 Haiku as judge for methodological consistency. Models sorted by SSI (descending).
Model Mean SSI $\uparrow$ Flip Rate % Unstable Refusal Rate
Gemma-3-12B-Instruct 0.965 18.4% 6.7% 78.5%
Llama-3.1-8B-Instruct 0.944 27.3% 10.4% 79.3%
Qwen3-8B 0.938 27.7% 11.8% 92.5%
Qwen2.5-7B-Instruct 0.938 26.3% 12.0% 81.3%
Figure 6 shows the response distribution (REFUSE, PARTIAL, COMPLY) for each model. The models differ in their base refusal rates: Qwen is more conservative (90.8% refuse) while Llama shows a slightly lower refusal rate (83.1%) with more PARTIAL responses (10.9% vs 3.9%).
@@ 5.1 Implications for Safety Evaluation
Deployment Considerations. For safety-critical applications, using deterministic sampling (temperature 0.0) provides the most stable behavior (mean SSI = 0.951). At this temperature, only 9.5% of prompts show instability, compared to 19.6% at temperature 1.0.
Adversarial Robustness. The existence of borderline prompts that flip between refuse and comply suggests that adversaries could exploit stochastic sampling by repeatedly querying the model until they obtain a harmful response.
@@ 5.3 Reproducibility Details
Inference Engine vLLM (latest)
@@
Repetition Penalty 1.0 (none)
@@
Judge Prompt See Appendix
@@ 6 Conclusion
Critically, single-shot evaluation—the standard practice in safety benchmarking—agrees with multi-sample ground truth only 92.4% of the time. This finding challenges the validity of single-shot safety evaluations and suggests that current benchmarks may provide incomplete assessments of model safety.

### v2

@@ Abstract
Current safety evaluations of large language models rely on single-shot testing, implicitly assuming that model responses are deterministic and representative of the model’s safety alignment. We challenge this assumption by investigating the stability of safety refusal decisions across random seeds and temperature settings. Testing four instruction-tuned models from three families (Llama 3.1 8B, Qwen 2.5 7B, Qwen 3 8B, Gemma 3 12B) on 876 harmful prompts across 20 different sampling configurations (4 temperatures $\times$ 5 random seeds), we find that 18–28% of prompts exhibit decision flips—the model refuses in some configurations but complies in others—depending on the model. Our Safety Stability Index (SSI) reveals that higher temperatures significantly reduce decision stability (Friedman $\chi^{2}=396.81$, $p<0.001$), with mean within-temperature SSI dropping from 0.977 at temperature 0.0 to 0.942 at temperature 1.0. We validate our findings across all model families using Claude 3.5 Haiku as a unified external judge, achieving 89.0% inter-judge agreement with our primary Llama 70B judge (Cohen’s $\kappa=0.62$). Within each model, prompts with higher compliance rates exhibit lower stability (Spearman $\rho=-0.47$ to 0.70-0.70, all $p<0.001$), indicating that models “waver” more on borderline requests. These findings demonstrate that single-shot safety evaluations are insufficient for reliable safety assessment and that evaluation protocols must account for stochastic variation in model behavior. We show that single-shot evaluation agrees with multi-sample ground truth only 92.4% of the time when pooling across temperatures (94.2–97.7% at fixed temperature depending on setting), and recommend using at least 3 samples per prompt for reliable safety assessment.
@@ 1 Introduction
  • Temperature significantly affects stability (Friedman $\chi^{2}=396.81$, $p<0.001$): Lower temperatures yield more stable decisions (mean within-temperature SSI = 0.977 at temp 0.0) compared to higher temperatures (mean SSI = 0.942 at temp 1.0)
@@ 3.1 Experimental Design
Dataset. We use 876 harmful prompts from the BeaverTails dataset [6], sourced from the 330k_train split filtered to is_safe=False, deduplicated, and taking the first 876 unique prompts. Categories include violence, hate speech, illegal activities, and other harmful content.
@@ 4.1 Overall Stability
The overall response distribution shows that the model predominantly refuses harmful requests (83.1% REFUSE, 10.9% PARTIAL, 5.9% COMPLY), suggesting strong safety alignment on average. However, the substantial proportion of unstable prompts indicates that aggregate statistics obscure significant per-prompt variability. Judge Note: Sections 4.1–4.3 report Llama 3.1 8B results using the Llama 70B judge (32% flip rate). Table 3 reports cross-model results using Claude 3.5 Haiku (27.3% flip rate for Llama). The 4.7 percentage point difference reflects inter-judge variation ($\kappa=0.62$); see Section 4.7 for validation.
@@ 4.2 Distribution of Stability Scores
Figure 2 presents the distribution of SSI scores across all 876 prompts. The distribution is heavily right-skewed with a strong mode at perfect stability (SSI = 1.0): the median SSI is 1.0, indicating that more than half of prompts receive identical classification labels across all configurations (though the response text may vary). However, 14.3% of prompts fall below the 0.8 stability threshold, demonstrating that instability affects a meaningful fraction of safety evaluations.
@@ 4.3 Temperature Effects
Figure 3 illustrates the relationship between temperature and both compliance rate and mean SSI. We observe that stability generally decreases as temperature increases (Friedman $\chi^{2}=396.81$, $p<0.001$, Kendall’s $W=0.038$ indicating a small but highly significant effect). Mean within-temperature SSI drops from 0.977 at temperature 0.0 to 0.942 at temperature 1.0, a decrease of 0.035 [95% CI: 0.032–0.039]. Post-hoc Wilcoxon signed-rank tests with Bonferroni correction confirm significant decreases in SSI between temperature 0.0 and 1.0 ($p<0.001$). The flip rate increases from 5.1% at temperature 0.0 to 23.6% at temperature 1.0 for Llama 3.1 8B (similar patterns across all models).
@@
Interestingly, the lowest compliance rate occurs at temperature 1.0 (3.0%). However, this may reflect that high-temperature sampling produces less coherent outputs that fail to meaningfully engage with the harmful request, rather than representing more robust safety behavior.
4.3.1 Within-Temperature Seed Stability
To separate seed variance from temperature-induced distribution shift, we computed SSI separately at each temperature using only the 5 seeds per configuration. Table 1 shows that even at temperature 0.0 (greedy decoding), there is non-zero instability: 5–12% of prompts flip across seeds depending on the model. This residual variation may arise from GPU non-determinism in floating-point operations. At temperature 1.0, flip rates increase to 12–24% depending on the model, confirming that higher temperatures substantially increase seed sensitivity.
Table 1: Within-temperature stability (seed variance only, N=5 seeds). Shows that instability increases with temperature even when holding temperature fixed.
Model t=0.0 t=0.3 t=0.7 t=1.0
Mean Within-Temperature SSI
Gemma-3-12B 0.986 0.972 0.967 0.964
Llama-3.1-8B 0.985 0.961 0.948 0.926
Qwen3-8B 0.967 0.948 0.943 0.945
Qwen2.5-7B 0.972 0.956 0.941 0.932
Flip Rate (%)
Gemma-3-12B 4.9 9.9 11.3 12.6
Llama-3.1-8B 5.1 12.7 17.2 23.6
Qwen3-8B 12.0 17.6 18.6 18.4
Qwen2.5-7B 9.2 15.4 20.1 22.8
@@ 4.5 What Predicts Prompt Instability?
To understand what makes certain prompts unstable, we conducted an exploratory analysis of linguistic features and their correlation with SSI. These analyses involve multiple comparisons and should be interpreted as hypothesis-generating rather than confirmatory. We find that some prompt characteristics correlate with instability:
  • Sensitive terms increase stability: Prompts containing obvious harmful keywords (e.g., “hack,” “bomb,” “kill”) have higher SSI (Pearson $r=0.18$), suggesting the model confidently refuses clearly harmful requests.
  • Multi-sentence prompts are less stable: Prompts with multiple sentences show lower SSI (Pearson r=0.21r=-0.21), possibly because complex requests create ambiguity.
A logistic regression classifier predicting instability from linguistic features achieves AUC = 0.73 [95% CI: 0.69–0.78] in 5-fold cross-validation, indicating that surface-level features partially explain which prompts will be unstable. [^1] Features used: word count, character count, average word length, sentence count, question mark presence, sensitive term count, code marker presence, and uppercase character ratio. The most predictive features are average word length, sensitive term count, and word count.
To quantify category effects, we classified prompts into 12 harm categories using keyword matching. Copyright-related requests (N=112) are dramatically more unstable than all other categories: mean SSI = 0.568 with 89.3% of prompts classified as unstable, compared to misinformation (N=30, SSI = 0.822, 20% unstable) and hacking (N=118, SSI = 0.862, 22% unstable). Copyright prompts also show the highest compliance rate (45.9%), suggesting the model is uncertain whether reproducing copyrighted content constitutes harm. In contrast, self-harm prompts (N=15) are most stable (SSI = 0.990, 0% unstable), likely because safety training provides clear guidance on this category. This suggests that instability concentrates in categories where harm boundaries are ambiguous. (Note: categories assigned via keyword matching; the “other” category contains 292 prompts not matching specific keywords.)
@@ 4.6 Model Comparison
Figure 5 compares stability across the four evaluated models. We observe consistent patterns of instability across all three model families, with instability rates (SSI << 0.8) ranging from 6.7% to 12.0%.
@@
Table 3: Stability comparison across model families with 95% bootstrap CIs. All models evaluated using Claude 3.5 Haiku as judge. Models sorted by SSI (descending).
Model Mean SSI $\uparrow$ Flip Rate % Unstable Refusal
Gemma-3-12B .965 [.958, .971] 18.3% [15.6, 20.9] 6.7% [5.3, 8.3] 78.6%
Llama-3.1-8B .944 [.936, .951] 27.3% [24.4, 30.1] 10.4% [8.4, 12.3] 79.2%
Qwen3-8B .938 [.929, .946] 27.8% [24.8, 30.9] 11.8% [9.6, 13.9] 92.5%
Qwen2.5-7B .938 [.929, .946] 26.3% [23.3, 29.1] 12.0% [9.9, 14.2] 81.3%
Figure 6 shows the response distribution (REFUSE, PARTIAL, COMPLY) for each model. The models differ substantially in their base refusal rates: Qwen 3 8B is most conservative (92.5% refuse), while Gemma 3 12B shows the lowest refusal rate (78.5%) with more PARTIAL responses (14.4% vs 6.5% for Qwen 3). Llama 3.1 8B and Qwen 2.5 7B fall in the middle (79.3% and 81.3% respectively).
Stability-Compliance Relationship. While the most stable model (Gemma) also has the lowest refusal rate, this correlation is based on only four models and may reflect model family differences. However, analyzing the relationship within each model provides stronger evidence: per-prompt SSI correlates negatively with compliance rate (Spearman $\rho=-0.47$ to 0.70-0.70 depending on model, all $p<0.001$). This indicates that prompts where the model is more likely to comply are also less stable—the model “wavers” more on borderline-harmful requests.
@@ 4.7 External Judge Validation
Bounding Judge Noise Contribution. To quantify how much measured instability reflects judge labeling ambiguity versus true model variation, we recomputed metrics using only responses where both judges agreed on classification. When restricting to consensus responses, mean SSI increased from 0.941 to 0.965, and flip rate decreased from 26.8% to 17.5%—a reduction of 9.3 percentage points. This 9.3pp reduction represents an upper bound on judge noise contribution: some excluded responses may reflect genuine model variation that happens to produce borderline outputs difficult for judges to classify consistently. Under this conservative interpretation, at least two-thirds of observed flip rate represents true model instability, with the remaining portion attributable to a combination of judge noise and genuinely ambiguous model outputs.
@@ 4.8 Sample Size Requirements for Reliable Evaluation
4.8.1 Fixed-Temperature Reliability
[…]
Table 6: Sample size reliability at fixed temperatures. Values show % agreement with 5-seed ground truth. At greedy decoding (t=0.0), single-shot evaluation is 97.7% reliable.
N Seeds t=0.0 t=0.3 t=0.7 t=1.0
1 97.7% 95.9% 95.0% 94.2%
2 97.6% 96.2% 95.2% 94.2%
3 98.9% 98.1% 97.7% 97.2%
5 100% 100% 100% 100%
4.9 Binary Scoring Ablation
[…]
Table 7: Binary scoring ablation: results under different PARTIAL handling rules. Instability persists regardless of scoring methodology.
Scoring Rule Mean SSI Flip Rate % Unstable
Three-way (original) 0.946 24.9% 10.2%
[…]
@@ 5.1 Implications for Safety Evaluation
Deployment Considerations. For safety-critical applications, using deterministic sampling (temperature 0.0) provides the most stable behavior (mean within-temperature SSI = 0.977). At this temperature, only 9.5% of prompts show instability, compared to 19.6% at temperature 1.0.
Adversarial Robustness. The existence of borderline prompts that flip between refuse and comply creates a practical vulnerability: adversaries can exploit stochastic sampling by repeatedly querying until obtaining a harmful response. We quantified this risk by computing expected queries to first compliance for prompts where any COMPLY response was observed (N=454 prompt-model pairs). The median expected queries is just 2, with 69% of exploitable prompts yielding compliance within 5 queries. This demonstrates that instability creates exploitable attack surfaces even for prompts that usually refuse.
@@ 5.2 Limitations
Temperature-Seed Pooling. Our primary SSI metric pools responses across both temperatures and seeds (N=20 configurations). Benchmarks that fix temperature may experience different reliability characteristics; we provide within-temperature analyses in Section 4.3.1 to address this. At greedy decoding (t=0.0), single-shot evaluation is 97.7% reliable, compared to 92.4% when pooling across temperatures.
Greedy Decoding Non-Determinism. We verified greedy decoding determinism by comparing response texts across seeds at temperature 0.0. Llama 3.1 8B showed 73.4% exact-match responses across seeds, while Qwen 2.5 7B showed only 17.6% exact matches, indicating substantial inference non-determinism even at temperature 0.0. This variation likely stems from GPU floating-point non-determinism in batched vLLM inference. Of prompts with byte-identical responses, judge label flips occurred in only 0.2% of cases for Llama, confirming that when responses are truly identical, judge labeling is highly consistent. The observed flip rates at t=0.0 (5–12%) thus reflect genuine inference variation rather than judge noise. To definitively separate inference artifacts from model behavior, future work should test with strict determinism settings (e.g., CUBLAS_WORKSPACE_CONFIG=:4096:8, torch.use_deterministic_algorithms(True)).
@@ 5.3 Reproducibility Details
Inference Engine vLLM 0.6.3
PyTorch Version 2.4.1
CUDA Version 12.1
GPU NVIDIA A100-80GB
Batch Size 32
@@
Determinism Flags None (default)
@@
Judge Prompt See below
@@ 6 Conclusion
Critically, single-shot evaluation—the standard practice in safety benchmarking—agrees with multi-sample ground truth only 92.4% of the time when pooling across temperatures (94.2–97.7% at fixed temperature depending on setting). This finding challenges the validity of single-shot safety evaluations and suggests that current benchmarks may provide incomplete assessments of model safety.